# KPC-cF: Aspect-Based Sentiment Analysis via Implicit-Feature Alignment with Corpus Filtering

### v6

@@ 7 Limitations and Future Work
Our analysis of the PL-cF thresholding reveals that while it reduces the number of predicted Target None classes resulting from extensive training on None class in TR, it also qualitatively explains why the error rate for Target None can be lower (see Table 7, 7). Contrary to experimental results that show it is reasonable to predict high-confidence Target Polarity labels when there is high surface similarity between sentences, cases where sentence similarity is less than 0.15 but predicted as Target None with confidence greater than 0.5 can hinder the model’s understanding of deeper semantic relationships. For instance, issues related to hygiene and cleanliness are not detected through the overall atmosphere, and in cases where multiple aspects overlap, the model fails to distinguish them clearly, incorrectly labeling them all as None. Additionally, aspects such as waiting space, waiting time, and waiting staff, where the boundary between service and atmosphere is ambiguous or falls into general anecdotes, are consistently mislabeled as None. Furthermore, our threshold filtering assigns a high-confidence Target None label for connotations with high similarity but lower errors, which supports a better distinction between semantic similarity and causal relationships. This is particularly effective when excluding personalized terms and Korean proper nouns, such as ‘Fixture (i.e., local native)’ used by reviewers. Our current translation-based training does not sufficiently capture the tone of sentences composed of proper nouns and particles from various domains within Korea. However, we have confirmed through various evaluations (Table 1, 2, 3, 4, 5, 7, 11 and Figure 5, 6, 7, 8) that our filtering approach aids in learning actual level of entailment relationships, moving beyond merely identifying similar relationships between sentences. To mitigate spurious correlations (Haig, 2003; Udomcharoenchaikit et al., 2022), we propose upper filtering that enables Target predictions to function as data facilitating the model’s learning of true entailment relationships between sentences, without relying solely on surface-level similarity in NLI pairs.
@@ Acknowledgements
We thank Chanseo Nam for manually correcting and annotating the Kor-SemEval and KR3 testset used to test our model, during his undergraduate research. We also thank Kyeongpil Kang, Byungok Han, Woohan Yun, Soyoung Yang, Joosun Yum and Guijin Son for their valuable discussions. Thanks to DAVIAN, MODULABS and UST AI as well. This research was supported by Brian Impact Foundation, a non-profit organization dedicated to the advancement of science and technology for all.
@@ A.1 Gradient Stability and Directionality from Dual Scoring
$|\text{sim}(x_{1},x_{2})-\text{sim}(x_{1},x_{3})|\leq L\cdot\|x_{2}-x_{3}\|_{2}$
After filtering, we observe:
@@
$\sigma_{C}^{f}$ $:=\sqrt{\mathbb{E}_{\mathcal{D}_{f}}[(C-\mu_{C}^{f})^{2}]}<\sqrt{\mathbb{E}_{\mathcal{D}}[(C-\mu_{C})^{2}]}=:\sigma_{C}$
Define filtered gradient:
@@
$\left\langle\nabla_{W_{H}}\ell(f_{t_{0}}(x_{s},x_{a}),y),\Phi(x_{s},x_{a})\right\rangle>0$
@@
$\left\langle\nabla_{W_{KQ}}\mathcal{L}_{f},\;W_{KQ}\Phi^{*}\right\rangle>0.$
@@
$P_{\text{post}}(y|x_{s},x_{a})-P_{\text{pre}}(y|x_{s},x_{a})>\delta>0$
indicating alignment with Assumption 3.2.
@@
Then for $L$-smooth $\mathcal{L}$ and Adam optimizer:
@@ D.2 Preliminary Results
Appendix E Attention Head Visualization

### v7

@@ 7 Limitations and Future Work
Our analysis of the PL-cF thresholding reveals that while it reduces the number of predicted Target None classes resulting from extensive training on None class in TR, it also qualitatively explains why the error rate for Target None can be lower (see Table 7, 7). Contrary to experimental results that show it is reasonable to predict high-confidence Target Polarity labels when there is high surface similarity between sentences, cases where sentence similarity is less than 0.15 but predicted as Target None with confidence greater than 0.5 can hinder the model’s understanding of deeper semantic relationships. For instance, issues related to hygiene and cleanliness are not detected through the overall atmosphere, and in cases where multiple aspects overlap, the model fails to distinguish them clearly, incorrectly labeling them all as None. Additionally, aspects such as waiting space, waiting time, and waiting staff, where the boundary between service and atmosphere is ambiguous or falls into general anecdotes, are consistently mislabeled as None. Furthermore, our threshold filtering assigns a high-confidence Target None label for connotations with high similarity but lower errors, which supports a better distinction between semantic similarity and causal relationships. This is particularly effective when excluding personalized terms and Korean proper nouns, such as ‘Fixture (i.e., local native)’ used by reviewers. Our current translation-based training does not sufficiently capture the tone of sentences composed of proper nouns and particles from various domains within Korea. However, we have confirmed through various evaluations (Table 1, 2, 3, 4, 5, 7, 12 and Figure 5, 6, 7, 8) that our filtering approach aids in learning actual level of entailment relationships, moving beyond merely identifying similar relationships between sentences. To mitigate spurious correlations (Haig, 2003; Udomcharoenchaikit et al., 2022), we propose upper filtering that enables Target predictions to function as data facilitating the model’s learning of true entailment relationships between sentences, without relying solely on surface-level similarity in NLI pairs.
@@ Acknowledgements
We thank Chanseo Nam for manually correcting and annotating the Kor-SemEval and KR3 testset used to test our model, during his undergraduate research. We also thank Kyeongpil Kang, Jinwuk Seok, Byungok Han, Woohan Yun, Guijin Son, Soyoung Yang and Joosun Yum for their valuable discussions. Thanks to DAVIAN, MODULABS and UST AI as well. This research was supported by Brian Impact Foundation, a non-profit organization dedicated to the advancement of science and technology for all.
@@ A.1 Gradient Stability and Directionality from Dual Scoring
$|\text{sim}(x_{1},x_{2})-\text{sim}(x_{1},x_{3})|\leq L\cdot\|x_{2}-x_{3}\|_{2}.$
After filtering, we observe in Table 2:
@@
$\sigma_{C}^{f}$ $:=\sqrt{\mathbb{E}_{\mathcal{D}_{f}}[(C-\mu_{C}^{f})^{2}]}<\sqrt{\mathbb{E}_{\mathcal{D}}[(C-\mu_{C})^{2}]}=:\sigma_{C}.$
Define filtered gradient such that:
@@
Proof of Lemma 1.
Suppose, towards a contradiction, that the gradient variance on $\mathcal{D}_{f}$ is not bounded by the stated $\varepsilon_{1}$. In particular, assume
$\mathbb{E}_{(x_{s},x_{a})\sim\mathcal{D}_{f}}\Big{[}\big{\|}\nabla_{\theta}\ell(f_{\theta}(x_{s},x_{a}),y)\;-\;\nabla_{\theta}L_{f}\big{\|}_{2}^{2}\Big{]}\;>\;\varepsilon_{1}(\sigma_{f}^{sim},\,\sigma_{f}^{C})\,.$
Consider any two samples $i,j\in\mathcal{D}_{f}$. By Lipschitz continuity of $sim$ and $C$ and the differentiability of $\ell$, there exist constants $K_{1},K_{2}$ such that the gradient difference is bounded by the score differences:
$\big{\|}\nabla_{\theta}\ell(f_{\theta}(x_{i}),y_{i})-\nabla_{\theta}\ell(f_{\theta}(x_{j}),y_{j})\big{\|}^{2}\;\leq\;K_{1}\,|\,sim_{i}-sim_{j}\,|^{2}\;+\;K_{2}\,|\,C_{i}-C_{j}\,|^{2}.$
Taking expectation over $i,j\sim\mathcal{D}_{f}$ gives an upper bound on the gradient variance:
$\mathrm{Var}_{\mathcal{D}_{f}}[\nabla_{\theta}\ell]\;\leq\;K_{1}\,(\sigma_{f}^{sim})^{2}\;+\;K_{2}\,(\sigma_{f}^{C})^{2}\;=\;\varepsilon_{1}(\sigma_{f}^{sim},\,\sigma_{f}^{C})\,.$
This contradicts the assumption that the variance exceeds $\varepsilon_{1}(\sigma_{f}^{sim},\sigma_{f}^{C})$. Hence, the gradient deviation on $\mathcal{D}_{f}$ is bounded by $\varepsilon_{1}(\sigma_{f}^{sim},\sigma_{f}^{C})$ as claimed. ∎
@@
$\left\langle\nabla_{W_{H}}\ell(f_{t_{0}}(x_{s},x_{a}),y),\Phi(x_{s},x_{a})\right\rangle>0,$
@@
Proof of Lemma 2.
Assume for contradiction that the initial gradient is not aligned with the fused representation; i.e. $\langle\,\nabla_{W_{H}}\ell(f_{t0}(x_{s},x_{a}),\,y),\;\Phi(x_{s},x_{a})\,\rangle\leq 0$ at the reinitialization step $t_{0}$. Let $p=p(y\mid x_{s},x_{a};\theta(t_{0}))$ be the model’s predicted probability for class $y$ at $t_{0}$. Since $W_{H}(t_{0})$ is random, $p<1$ for the true class $y$. The gradient of the loss w.r.t. the classification weight for class $y$ is
$\nabla_{W_{H},y}\,\ell(f_{t0}(x_{s},x_{a}),y)\;=\;\big{(}p-1\big{)}\,\Phi(x_{s},x_{a})\,,$
by the derivative of cross-entropy. The gradient descent update is $-\,\nabla_{W_{H},y}\ell=(1-p)\,\Phi(x_{s},x_{a})$. Its inner product with $\Phi(x_{s},x_{a})$ is
[…]
since $01-p>$ 0. This means the weight update at $t_{0}$ is positively aligned with $\Phi(x_{s},x_{a})$, i.e. $\langle\nabla_{W_{H}}\ell,\;\Phi\rangle>0$. This contradicts our assumption that $\langle\nabla_{W_{H}}\ell,\;\Phi\rangle\leq 0$. Therefore, the initial gradient is indeed aligned with the semantic direction $\Phi(x_{s},x_{a})$, yielding $\langle\nabla_{W_{H}}\ell(f_{t0}(x_{s},x_{a}),y),\;\Phi(x_{s},x_{a})\rangle>0$. ∎
Tables 1 and 4 show that, for semantic label detection, filtered data consistently yields higher ACP scores than unfiltered data. Even the BaselinemBERT+TR+PL-cF model—which exhibits a relatively low transfer learning effect in Table 1—achieves higher ACP with filtered data. Likewise, Baseline+PL variants without TR training also show improved ACP under filtering, further supporting our lemma.
@@
$\left\langle\nabla_{W_{KQ}}\mathcal{L}_{f},\;W_{KQ}\Phi^{*}\right\rangle>0,$
indicating alignment with Assumption 3.1, Figure 6, 7, 8, Table 3 and enhanced ACD score in Table 1.
@@
$P_{\text{post}}(y|x_{s},x_{a})-P_{\text{pre}}(y|x_{s},x_{a})>\delta>0,$
indicating alignment with Assumption 3.2 and Figure 3.
Proof of Lemma 4.
Assume the fine-tuning yields no increase in the predicted probability for the true class $y$. Let $p_{\text{pre}}=P_{\text{pre}}(y\mid x_{s},x_{a})$ be the model’s probability for $y$ with the initial classifier $W_{H}(t_{0})$, and $p_{\text{post}}=P_{\text{post}}(y\mid x_{s},x_{a})$ after training. Our assumption is $p_{\text{post}}\leq p_{\text{pre}}$. However, from Lemma 2 we know the initial gradient update increases the logit for class $y$. In particular, after one gradient step of size $\eta$, the logit $z_{y}$ for class $y$ satisfies
$z_{y}(t_{0}+\!1)\;=\;z_{y}(t_{0})\;+\;\eta\,(1-p_{\text{pre}})\,\|\Phi(x_{s},x_{a})\|^{2},$
so $z_{y}(t_{0}+1)>z_{y}(t_{0})$ since $01-p_{\text{pre}}>$ 0. This implies $P(y\mid x_{s},x_{a})$ increases after the first update: $p^{(1)}:=P(y|x_{s},x_{a})_{t_{0}+1}>p_{\text{pre}}$. By local convexity of $\ell$ (Assumption 3), further gradient steps will continue to improve the prediction on this sample. Therefore $p_{\text{post}}\geq p^{(1)}>p_{\text{pre}}$, contradicting the assumption $p_{\text{post}}\leq p_{\text{pre}}$. We conclude $p_{\text{post}}-p_{\text{pre}}>0$. In fact, one can take $\delta:=p^{(1)}-p_{\text{pre}}>0$ as a guaranteed improvement, so $P_{\text{post}}(y|x_{s},x_{a})-P_{\text{pre}}(y|x_{s},x_{a})>\delta$. ∎
@@
Then for $L$-smooth $\mathcal{L}$ and AdamW optimizer:
@@
Proof of Convergence Theorem.
Assume, to reach a contradiction, that $\|\nabla_{\theta}L(\theta_{t})\|_{2}$ does not converge to 00. Then there exists some $\epsilon>0$ and an infinite sequence of iterations $\{t_{k}\}$ such that $\|\nabla_{\theta}L(\theta_{t_{k}})\|_{2}\geq\epsilon$ for all $k$. Under the bounded gradient dispersion assumption, the Adam optimizer’s adaptive step size can be treated as a positive constant $\eta>0$ (for sufficiently large $t_{k}$). Because $L(\theta)$ is $L$-smooth, the descent lemma gives for a small enough $\eta$ (e.g. $\eta\leq 1/L$):
$L(\theta_{t+1})\;\leq\;L(\theta_{t})\;-\;\eta\,\Big{(}1-\frac{L\,\eta}{2}\Big{)}\,\|\nabla_{\theta}L(\theta_{t})\|_{2}^{2}.$
Whenever $\|\nabla_{\theta}L(\theta_{t})\|_{2}\geq\epsilon$, this implies $L(\theta_{t+1})\leq L(\theta_{t})-c\,\epsilon^{2}$ for some constant $c=\eta(1-\tfrac{L\eta}{2})>0$. In other words, every iteration with large gradient norm yields at least a fixed reduction $c\,\epsilon^{2}$ in the loss. After $N$ such iterations (with $\|\nabla L\|\geq\epsilon$), one would have
$L(\theta_{t_{N}})\;\leq\;L(\theta_{t_{0}})\;-\;N\,c\,\epsilon^{2}.$
For large $N$, the right-hand side becomes negative, which is impossible since $L(\theta)$ (being an average cross-entropy) is bounded below by 00. This contradiction implies our assumption was false. Therefore $\|\nabla_{\theta}L(\theta_{t})\|_{2}\to 0$ as $t\to\infty$, i.e. the training converges to a stationary point. ∎
@@
Preliminaries and Assumptions
Setting. Let $\Phi(x_{s},x_{a})\in\mathbb{R}^{d}$ be a token-embedding feature for inputs $(x_{s},x_{a})$. We assume $\Phi$ maps into some subset $\mathcal{E}\subset\mathbb{R}^{d}$.
Assumption (1): $\mathcal{E}$ is Compact
Since $\Phi(x_{s},x_{a})\in\mathcal{E}$ for all $(x_{s},x_{a})$, we want $\mathcal{E}$ to be a compact subset of $\mathbb{R}^{d}$. By the Heine–Borel theorem, compactness in $\mathbb{R}^{d}$ is equivalent to being closed and bounded. Hence:
  1. 1.
    Boundedness. There exists $R>0$ such that
    $\|\Phi(x_{s},x_{a})\|_{2}\;\leq\;R\quad\text{for all }(x_{s},x_{a}).$
  2. 2.
    Closedness. For any convergent sequence $\{\Phi(x_{s}^{n},x_{a}^{n})\}_{n=1}^{\infty}\subset\mathcal{E}$, its limit also lies in $\mathcal{E}$. Equivalently, $\mathcal{E}$ contains all its limit points.
Assumption (2): $\mathrm{sim}$ and $C$ are Lipschitz-Continuous
Define
$\mathrm{sim}(x_{s},x_{a})\;=\;\frac{\langle x_{s},x_{a}\rangle}{\|x_{s}\|\cdot\|x_{a}\|},\quad C(x_{s},x_{a})\;=\;\max_{k}\,p\bigl{(}k\mid x_{s},x_{a};\theta\bigr{)}.$
Each must be Lipschitz in $\Phi(x_{s},x_{a})$. Formally, there is an $L_{f}>0$ such that
$\bigl{|}\,f(x_{s},x_{a})-f(x_{s}^{\prime},x_{a}^{\prime})\bigr{|}\;\leq\;L_{f}\,\|\Phi(x_{s},x_{a})-\Phi(x_{s}^{\prime},x_{a}^{\prime})\|_{2},\quad\forall(x_{s},x_{a}),(x_{s}^{\prime},x_{a}^{\prime})\in\mathcal{D}.$
Assumption (3): $\ell$ is Locally Convex and Differentiable
Let $\ell\bigl{(}f_{\theta}(\Phi(x_{s},x_{a})),y\bigr{)}$ be a loss (e.g. cross-entropy). We require $\ell$ to be differentiable, and in each local neighborhood it behaves like a convex function, ensuring standard gradient-based arguments near a minimum.
@@ C.1 Hyperparameter
Hyperparameter Value
Model xlm-roberta-base, bert-base-multilingual-cased
Learning rate 2e-5
Optimizer AdamW
Weight decay 0.01
Dropout (hidden layers & attention) 0.1
Hidden size 768
Intermediate size 3072
Vocabulary size 250K, 110K
Table 9: Hyperparameters of main phase for post-training on KR3 dataset
@@ D.2 Preliminary Results
Appendix E Attention-Head Visualization